OpenStudy (anonymous):

simplify the expression (5+6i)/ (4+6i)

OpenStudy (anonymous):

please help someone!

OpenStudy (anonymous):

Is that every bit of the problem?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@incomplte ??

OpenStudy (anonymous):

@jim_thompson5910 ??

OpenStudy (mertsj):

You need to multiply the numerator and denominator by the conjugate of the denominator which is 4-6i

OpenStudy (anonymous):

im still kinda confused @Mertsj

OpenStudy (anonymous):

would it be 56-6i/ 52

OpenStudy (mertsj):

\[\frac{5+6i}{4+6i}\times\frac{4-6i}{4-6i}\]

OpenStudy (mertsj):

The denominator is right.

OpenStudy (mertsj):

So is the numerator. Good job!!

OpenStudy (anonymous):

56-54i/ 52 is this it

OpenStudy (anonymous):

@Mertsj ??

OpenStudy (mertsj):

yes?

OpenStudy (anonymous):

what answer was correct

OpenStudy (mertsj):

Did you read where I told you the numerator is also correct?

OpenStudy (mertsj):

56 lala2 Best Response 056-54i/ 52 is this it